SpaceRef

SpaceRef


NASA/ESA SOHO Operations Day - 170/2003 (19 June 03)

Status Report From: Goddard Space Flight Center
Posted: Monday, June 23, 2003

SOHO DAILY REPORT

Operations Day - 170/2003 (19 June 03)

Mission Day 2757

SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC

Ref. #: SH-FOT-DR-030619

  • PASS #1 Madrid DSS66 0549-1713
  • PASS #2 Goldstone DSS16 1703-2140
  • PASS #3 Canberra DSS46 2118-171/0445

PASS #1 Madrid DSS66 0549-1713

Anomalies:

None

NRT:

1225-1257 1341-1355 1527-1547

Activities Summary:

  • - SSR Dump
  • - Collected MDI-M (0653-0710, 1304-1701)
  • - Collected MDI-H (0710-1304, 1701-1713)
  • - Reaction Wheel Speed Monitoring Daily Adjustment
  • - CRS Drift Estimation Update
  • - Hx Limit Update
  • - CAE-A in the AOCS HW Config Word
  • - SSR SEF Reset (x1)
  • - Executed OCD #1791: Initial HGA Unstick Activities
  • - Executed OCD #1792: Switch Off Red. APMM Heater/Restore Nom. Heater Settings

PASS #2 Goldstone DSS16 1703-2140

Anomalies:

None

NRT:

None

Activities Summary:

  • - Collected MDI-M (1741-2115)
  • - Collected MDI-H (1703-1741, 2115-2140)
  • - Executed OCD #1793: Use Both Nominal and Redundant APME Simultaneously

PASS #3 Canberra DSS46 2118-171/0445

Anomalies:

None

NRT:

2153-0038

Activities Summary:

  • - Collected MDI-M (2151-0321)
  • - Collected MDI-H (2118-2138, 2150-2151, 0321-0406, 0412-0445)
  • - OBT Distribution to SWAN, CELIAS
  • - Collected Ranging Data
  • - SSR SEF Reset (x2)

-- End of Report --

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

NanoRacks